OpenStudy (anonymous):

Evaluate the expression for the given value of the variable. 3(8 - x) - 2 for x = 0, 1, 2, and 3.

OpenStudy (kkutie7):

3(8-0)-2= 3(8)-2=24-2=22 3(8-1)-2= 3(7)-2=21-2=19 3(8-2)-2= 3(6)-2=18-2=16 3(8-3)-2= 3(5)-2=15-2=13